OpenStudy (anonymous):

A quadratic equation has an x-squared term, an x term and a constant term. Sometimes, Never, or Always?

OpenStudy (anonymous):

I think the answer is always but i'm not sure??

OpenStudy (anonymous):

see for a quadratic equation x^2 is must.......

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Always: Quadratic equation is shown by: \[\large ax^2 + bx + c = 0\]

OpenStudy (anonymous):

but x^2+3x=0.....what about this???

OpenStudy (anonymous):

Is it not quadratic??

OpenStudy (anonymous):

Yes it is here the constant term is 0..

OpenStudy (anonymous):

\[\large x^2 + 3x +0 = 0\]